OpenStudy (anonymous):

absolute value value equations |3+7x|+4=22

jigglypuff314 (jigglypuff314):

isolate the absolute values by subtracting 4 from both sides then know it can be separated into two equations like |a+b|=c to a+b=c and a+b = -c

OpenStudy (anonymous):

can you like set the problem up for me please that will help alot and i can do the rest @jigglypuff314

jigglypuff314 (jigglypuff314):

|3+7x|+4=22 -4 -4 |3+7x| = 18 so 3 + 7x = 18 and 3 + 7x = -18 so solve for x for both

OpenStudy (anonymous):

thx @jigglypuff314